OpenStudy (anonymous):

Use the geometric mean to find the 7th term in a geometric sequence if the 6th term is 50 and the 8th term is 18.

OpenStudy (anonymous):

A geometric sequence means you're multiplying by the same number over and over again to get each term in the sequence. Let's call that number x. So to get from the 6th to the 7th, you multiply the 6th by x, and to get from the 7th to the 8th, you multiply by x again. This information is enough to set up an equation to solve for x. Can you give it a shot?

OpenStudy (anonymous):

I just dont get how I'm supposed to start if I don't know what x is

OpenStudy (anonymous):

You can set up an equation. You take 50 and multiply it by x, and then multiply it by x again, and now you're at 18. What would that look like as an equation, and how can you solve for x?

OpenStudy (anonymous):

50x^2=18?

OpenStudy (anonymous):

would I divide 50 by 18?

OpenStudy (anonymous):

no, 18 by 50

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Yes. If you simplify the fraction, this will be a much easier problem.

OpenStudy (anonymous):

I got .36 after dividing 18 by 50.

OpenStudy (anonymous):

simplified it's 9/25

OpenStudy (anonymous):

Okay, so don't forget about the x^2...you're not done solving for x quite yet.

OpenStudy (anonymous):

I'm so confused. I don't know where to put anything

OpenStudy (anonymous):

Here's what you've done so far. 50x^2 = 18 x^2 = 18/50 x^2 = 9/25 Now how do you get rid of the "^2"?

OpenStudy (anonymous):

take the square root?

OpenStudy (anonymous):

Yes. And when you take the square root of a fraction, you just square root the top and square root the bottom. So what do you get for x?

OpenStudy (anonymous):

3/5?

OpenStudy (anonymous):

Yes! So now you have the number that you need to multiply by each time to get every term in the sequence: 3/5. So if the 6th term is 50, what is the 7th term?

OpenStudy (anonymous):

30. Thank you so much!

OpenStudy (anonymous):

Good job. You're welcome.